OpenStudy (anonymous):

Evalute 2^-3-4^-2 need to somehow get 1/16

OpenStudy (anonymous):

2^-3-4^-2 2^-3 = 1/(2*2*2) = 1/8 = 2/16 4^-2 = 1/(4*4) = 1/16 2/16-1/16 = (2-1)/16 = 1/16 done.